OpenStudy (miguel980):

Bob is 7 years older than Mary. Five years from now, he will be twice her age. Find both of their ages

OpenStudy (jango_in_dtown):

Let the present age of Mary be x years. Then present age of Bob=7+x By question (7+x)+5=2x implies x=12 Therefore Mary is 12 years old and bob is 7+12=19 years old